OpenStudy (icecreamyummy):

will fan and may be metal Mia deposits either $5 or $3.50 into her savings account. She made 14 deposits equaling a total of $59.50. How many times did she deposit $5 and how many times did she deposit $3.50? Mia made deposits of $5 and deposits of $3.50

OpenStudy (icecreamyummy):

@gottennis121 @iambatman @HelpBlahBlahBlah @chycora_thebaddest14 @Armoore

OpenStudy (icecreamyummy):

@demonchild99 I know you good with math

OpenStudy (demonchild99):

divide the amt of 59.50 by 5 then 59.50 bye 3.50

OpenStudy (demonchild99):

and there is your answers

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

hi

OpenStudy (misty1212):

lets call \(x\) the amount of $5 deposits ok?

OpenStudy (icecreamyummy):

k

OpenStudy (misty1212):

then since the total is 14, the amount of $3.50 deposits is \(14-x\)

OpenStudy (misty1212):

we have \[5x+3.50(14-x)\] as the total amount of money so we can solve \[5x+3.50(14-x)=59.50\]

OpenStudy (gottennis121):

are you doing the total? not each one separate into 59.50?

OpenStudy (misty1212):

may be easier if we multiply both sides by ten and solve \[50x+35(14-x)=595\]

OpenStudy (misty1212):

multiply out and combine like terms on the left to get \[50x+490-35x=595\] or \[15x+490=595\]

OpenStudy (misty1212):

subtract 490 from both sides, then divide by 15 to get your answer !!

OpenStudy (icecreamyummy):

x=7

OpenStudy (icecreamyummy):

but I have two blanks to fill

OpenStudy (icecreamyummy):

@misty1212

OpenStudy (misty1212):

yes you do!

OpenStudy (misty1212):

what was x? the number of $5 deposits or the number of 3.50 ones?

OpenStudy (icecreamyummy):

7 $5

OpenStudy (icecreamyummy):

@misty1212 is that rigth

OpenStudy (misty1212):

yes

OpenStudy (misty1212):

and the total number of deposits was 14, so how many are left?

OpenStudy (icecreamyummy):

7 o okay 7 fives and 7 3.50

OpenStudy (icecreamyummy):

cool thx and have a metal

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]

OpenStudy (misty1212):

thanks have a \(\color\red\heartsuit\)

OpenStudy (icecreamyummy):

thx